OpenStudy (anonymous):

Please find the exact value of the radical expression in simplest form. r√r - 3r√r + 4r√r

OpenStudy (kingj):

since they are all the square root of r you can add them like adding integers

OpenStudy (kingj):

1-3+4 = -2+4 2 answer is 2 square root of r

OpenStudy (anonymous):

Tell me what is: 1 - 3 + 4 = ??